### v1

@@ Abstract
Context:The existence of low-mass giants with large amounts of lithium (Li) in their surfaces has challenged stellar evolution for decades. One of the possibilities usually discussed in the literature to explain these Li-rich giants involves the interaction with a close binary companion, a scenario that predicts that, when compared against their non-enriched counterparts, Li-rich giants should preferentially be found as part of binary systems.
Aims:We aim to assemble the largest possible sample of low-mass giants with well-measured Li abundances, to determine with high statistical significance the close binary fractions of Li-rich and Li-normal giants, and thus test the binary interaction scenario for the emergence of Li-rich giants.
Methods:We develop a method that uses radial velocities (RVs) at three different epochs to quantify the degree of RV variability, which we use as a proxy for the presence of a close binary companion. The method is tested and calibrated against samples of known RV standard stars and known spectroscopic binaries. We then assemble a sample of 1418 giants with available RVs from RAVE, GALAH, and Gaia, as well as stellar parameters and Li abundances from GALAH, to which we apply our variability classification. We can determine an evolutionary state for 1030 of these giants. We also compare the results of our RV variability analysis with binarity indicators from the Gaia mission.
Results:When applying our methodology to the control samples, we find that the accuracy of the classification is controlled by the precision of the RVs used in the analysis. For the set of RVs available for the giants, this accuracy is 80-85%. Consistent with seismic studies, the resulting sample of giants contains a fraction of Li-rich objects in the red clump (RC) that is twice as large as that in the first ascent red giant branch (RGB). Among RC giants, the fractions of Li-rich objects with high RV variability and with no RV variability are the same as those for Li-normal objects, but we find some evidence that these fractions may be different for giants in the first-ascent RGB. Analysis of binary indicators in Gaia DR3 shows a smaller fraction of binary giants than our criteria, but no relation can be seen between Li enrichment and binarity either.
Conclusions:Our RV variability analysis indicates that there is no preference for Li-rich giants in the RC to be part of binary systems, thus arguing against a binary interaction scenario for the genesis of the bulk of Li-rich giants at that evolutionary stage. On the other hand, Li-rich giants in the RGB appear to have a small but measurable preference for having close companions, something that deserves further scrutiny with more and better data. Additional measurements of the RVs of these giants at higher RV precision would greatly help in confirming and more robustly quantifying these results.
Are Lithium-Rich Giants Binaries?: A Radial Velocity Variability Analysis of 1,400 Giants
Matias ${}^{,}$ 1122    Julio Chanamé 11       Claudia Aguilera-Gómez 11
(Received XXXX ; accepted YYYY)
@@ 1 Introduction
The source of Li in these anomalous giants remains a puzzle for stellar evolution. Diverse mechanisms have been invoked to explain the high Li abundance of giants. Some of these require an external source for the Li, such as the engulfment of planets (Siess & Livio 1999) or mass transfer from an asymptotic giant branch star that can produce Li by hot bottom burning (Sackmann & Boothroyd 1992). Some other hypotheses consider Li production in the interior of the star through the Cameron-Fowler mechanism (Cameron & Fowler 1971), requiring the rapid transport of elements between the region of nuclear burning, where fresh ${}^{7}$e or Li are produced, and the surface of the star. Depending on the efficient extra-mixing mechanism considered, and when this is triggered it could explain Li-rich giants at different evolutionary phases, such as the upper RGB or RC (e.g. Schwab 2020; Denissenkov et al. 2023). Additionally, some processes require the presence of an external source to trigger the internal production of Li, either the tidal interaction of a binary companion (Casey et al. 2019), or the merger with a helium-core white dwarf (Zhang et al. 2020). The Cameron-Fowler scenario for Li enhancement needs a fast and efficient mixing mechanism to increase the Li abundance of the giant. Casey et al. (2019) suggest that differential rotation that has been enhanced by the presence of a binary companion could produce the efficient levels of internal transport needed for Li enrichment. The additional angular momentum and tidal interaction produced by this companion would not only increase the surface Li but also the surface rotation rate. An interesting prediction of this model is that most of the Li-rich giant stars that are in the core-helium burning stage should have a binary companion since the sudden contraction of the giant at the beginning of the helium burning further enhances the transport of angular momentum in the stellar interior. It is therefore of large interest to determine the incidence of binaries among Li-rich giants.
@@ 2.2 Calibration against close binaries and RV stds
We find that $96.4\%$ of stars in the Cluster 0 were Class $\leq 4$ ($100\%$ for all subsets) according to our threshold for the $w$-parameter space, while $100\%$ of stars in the Cluster 1 were Class 5a or 5b ($93.7\%$ for all subsets). Therefore, the clustering algorithm naturally reproduces with high correspondence the separation between RV variable and non-variable stars achieved by the criteria in Table 1. This validates the use of those criteria for our purposes of identifying binaries among giants with Li measurements.
@@ 3 The RV variability of giants and their Li content
The sample of giants to be analyzed is constructed as follows. We first took all stars in Gaia DR3 with RV determinations, and crossmatched them with GALAH DR3 (so that Li measurements are available) and RAVE DR6, to produce a sample of stars with three independent RV determinations. Starting from this sample, we then selected giants following the recipe by Martell et al. (2021) (their Section 2.2), which uses the stellar parameters from GALAH. Then, we took those stars with no issues in their data, reduction, analysis, and iron abundance determination, summarized by the following flags in GALAH DR3: flag_sp == 0, flag_fe_h == 0, extinction parameter E(B-V)¡0.33, and photometric quality flagged as A, B, or C in the WISE $W_{2}$ band (added in the catalog of GALAH) which is later used for classifying the evolutionary phase of the giants (see Buder et al. 2021, for further details of the parameters). Secondly, the sample is restricted to a range of effective temperature $3000$ K $\leq T_{\mathrm{eff}}\leq$ $5730$ K and surface gravity log(g)$\leq 3.2$. Next, we identified if there were any stars belonging to the SMC or LMC by using their spatial and kinematic properties as described by Martell et al. (2021), but no members were found in our sample.
@@
We found a total of $5667$ red giants which have RV measurements with their respective errors reported by the three surveys. We also separately selected those stars that have flag_li_fe == 0 in GALAH DR3, which indicates a line detection good enough to consider a Li abundance measurement without upper limits and flagged problems (Martell et al. 2021; Buder et al. 2021). This subset contains a total of $1418$ red giants. Li abundances were obtained using the logarithmic number density defined as A(X)=$\log(N_{X}/N_{H})+12$ for an element X. From the abundances available in GALAH, i.e., [Li/Fe] and [Fe/H], we computed A(Li)=[[[[Li/Fe]]]]+[[[[Fe/H]]]]+1.05, where ${}_{\odot}$=1.05 is the abundance value of the Sun from Asplund et al. (2009). The errors for [Li/Fe] and [Fe/H] are given in GALAH DR3, then the errors for the Li abundance are obtained as eA(Li)=$\sqrt{\mathrm{e[Li/Fe]}^{2}+\mathrm{e[Fe/H]}^{2}}$. We found that the mean error is eA(Li)$\sim 0.1$ and the maximum is eA(Li)$\sim 0.25$. Since the A(Li) distributions are later analyzed with a bin size larger than the maximum error (see Section 3.1), we do not apply any additional selection based on eA(Li).
@@ 3.1 RV variability and Li abundance
The right-hand panel of Figure 6 zooms into the high-end tail of the A(Li) distributions from the left-hand panel, starting from A(Li)=1.5. In this panel, we consider joint classes. Thus, $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a joint class for a given range of A(Li). Giants in class 5a-5b are therefore those Li-rich giants with high certainty of being in a binary system (because they have the largest variation in RV), and they correspond to $37.6\%$ of all the Li-rich stars in our sample. This is only about $5\%$ larger than the fraction of stars with high certainty of variability when considering all giants with A(Li) measurement ($32.9\%$) and all giants with RV in the three surveys ($32.3\%$). Considering error bars, $33\%$ is about one standard deviation from the fraction of Li-rich giants in class 5a-5b (37.6%) which is marginal. Therefore, regardless of their Li abundance, there are about one-third of giants with a high probability of being in a binary system. At the same time, among the Li-rich giants only, there are no significant differences among the A(Li) distributions, they all show that the fraction $\mathrm{N_{s}/N_{class}}$ of giants in each A(Li) bin is almost the same for all the variability classes.

### v2

@@ Abstract
Context. The existence of low-mass giants with large amounts of lithium (Li) in their surfaces has challenged stellar evolution for decades. One of the possibilities usually discussed in the literature to explain these Li-rich giants involves the interaction with a close binary companion, a scenario that predicts that, when compared against their non-enriched counterparts, Li-rich giants should preferentially be found as part of binary systems.
Aims. We aim to assemble the largest possible sample of low-mass giants with well-measured Li abundances, to determine with high statistical significance the close binary fractions of Li-rich and Li-normal giants, and thus test the binary interaction scenario for the emergence of Li-rich giants.
Methods. We develop a method that uses radial velocities (RVs) at three different epochs to quantify the degree of RV variability, which we use as a proxy for the presence of a close binary companion. The method is tested and calibrated against samples of known RV standard stars and known spectroscopic binaries. We then assemble a sample of 1418 giants with available RVs from RAVE, GALAH, and Gaia, as well as stellar parameters and Li abundances from GALAH, to which we apply our variability classification. We can determine an evolutionary state for 1030 of these giants. We also compare the results of our RV variability analysis with binarity indicators from the Gaia mission.
Results. When applying our methodology to the control samples, we find that the accuracy of the classification is controlled by the precision of the RVs used in the analysis. For the set of RVs available for the giants, this accuracy is 80-85%. Consistent with seismic studies, the resulting sample of giants contains a fraction of Li-rich objects in the red clump (RC) that is twice as large as that in the first ascent red giant branch (RGB). Among RC giants, the fractions of Li-rich objects with high RV variability and with no RV variability are the same as those for Li-normal objects, but we find some evidence that these fractions may be different for giants in the first-ascent RGB. Analysis of binary indicators in Gaia DR3 shows a smaller fraction of binary giants than our criteria, but no relation can be seen between Li enrichment and binarity either.
Conclusions. Our RV variability analysis indicates that there is no preference for Li-rich giants in the RC to be part of binary systems, thus arguing against a binary interaction scenario for the genesis of the bulk of Li-rich giants at that evolutionary stage. On the other hand, Li-rich giants in the RGB appear to have a small but measurable preference for having close companions, something that deserves further scrutiny with more and better data. Additional measurements of the RVs of these giants at higher RV precision would greatly help in confirming and more robustly quantifying these results.
@@ 1 Introduction
The source of Li in these anomalous giants remains a puzzle for stellar evolution. Diverse mechanisms have been invoked to explain the high Li abundance of giants. Some of these require an external source for the Li, such as the engulfment of planets (Siess & Livio 1999) or mass transfer from an asymptotic giant branch star that can produce Li by hot bottom burning (Sackmann & Boothroyd 1992). Some other hypotheses consider Li production in the interior of the star through the Cameron-Fowler mechanism (Cameron & Fowler 1971), requiring the rapid transport of elements between the region of nuclear burning, where fresh 7Be or Li are produced, and the surface of the star. Depending on the efficient extra-mixing mechanism considered, and when this is triggered it could explain Li-rich giants at different evolutionary phases, such as the upper RGB or RC (e.g. Schwab 2020; Denissenkov et al. 2023). Additionally, some processes require the presence of an external source to trigger the internal production of Li, either the tidal interaction of a binary companion (Casey et al. 2019), or the merger with a helium-core white dwarf (Zhang et al. 2020). The Cameron-Fowler scenario for Li enhancement needs a fast and efficient mixing mechanism to increase the Li abundance of the giant. Casey et al. (2019) suggest that differential rotation that has been enhanced by the presence of a binary companion could produce the efficient levels of internal transport needed for Li enrichment. The additional angular momentum and tidal interaction produced by this companion would not only increase the surface Li but also the surface rotation rate. An interesting prediction of this model is that most of the Li-rich giant stars that are in the core-helium burning stage should have a binary companion since the sudden contraction of the giant at the beginning of the helium burning further enhances the transport of angular momentum in the stellar interior. It is therefore of large interest to determine the incidence of binaries among Li-rich giants.
@@ 2.1 RV variability as a binary indicator
To construct the variability classifications in Table 1, we conducted a qualitative assessment of our set of giants and grouped them in different subsamples, considering the observable differences in the RV time series and the blending of error bars at arbitrary epochs (as shown in the right-hand subpanels of Figure 1). By doing so, we defined distinct subsamples based on observable characteristics. We established representative thresholds in $w_{12}$ and $w_{t}$ that mostly enclose the features observed in each group of variability and we made sure that these thresholds described non-overlapping zones in the $w_{12}$-$w_{t}$ while encompassing the entire parameter space.
@@ 2.2 Calibration against close binaries and RV stds
We find that $96.4\%$ of stars in Cluster 0 were Class $\leq 4$ ($100\%$ for all subsets) according to our threshold for the $w$-parameter space, while $100\%$ of stars in Cluster 1 were Class 5a or 5b ($93.7\%$ for all subsets). Therefore, the clustering algorithm naturally reproduces with high correspondence the separation between RV variable and non-variable stars achieved by the criteria in Table 1. This validates the use of those criteria for our purposes of identifying binaries among giants with Li measurements. It is important to note that we consider three different classification groups: Those determined by the clustering algorithm (Cluster 0 or 1), those identified through extensive RV previous studies (SB9 or RV standards), and those classified based on the variability classes outlined in Section 2.1. While the clustering algorithm shows a separation similar to our threshold for high variability (Class 5a-b vs Class $\leq 4$), not all the objects categorized in Cluster 0 and 1 necessarily correspond to SB9 and RV standards, respectively. This occurs because our variability classification inherently involves some level of error when identifying binaries and RV standards, as we show in the top panel of Figure 3.
@@ 3 The RV variability of giants and their Li content
As previously discussed, we compiled RV measurements from Gaia DR3, GALAH DR3, and RAVE DR6. The Gaia mission uses the RVS spectrograph with a resolving power of $R=\lambda/\Delta\lambda\sim 11500$ for near-infrared spectra. The Gaia DR3 median formal precision for the RV data is of the order of $\sim 1\ \mathrm{km\ s^{-1}}$ (Katz et al. 2022). GALAH uses the HERMES spectrograph at the Australian Astronomical Observatory, reaching a resolution of $R\sim 28000$ in the optical wavelength range. The RV uncertainties in the GALAH DR3 survey are typically $\lesssim 0.1\ \mathrm{km\ s^{-1}}$ (Zwitter et al. 2021). RAVE uses the 6dF multi-object spectrograph at the UK Schmidt telescope at Siding Spring in Australia, corresponding to a resolving power of $R\sim 7500$ for a wavelength range of 8410–8795 Å. The RV uncertainties in RAVE DR6 are $<2\ \mathrm{km\ s^{-1}}$ (Steinmetz et al. 2020). Gaia and GALAH report barycentric RV measurements, whereas RAVE provides heliocentric data. Since the difference between both frames is $\lesssim 15\ \mathrm{m\ s^{-1}}$ (e.g. Nidever et al. 2002; Collier Cameron et al. 2021), much less than the typical uncertainties of the RV measurements in the surveys, we ignore any correction related to the reference of the RV data.
@@
The sample of giants to be analyzed is constructed as follows. We first took all stars in Gaia DR3 with RV determinations, and crossmatched them with GALAH DR3 (so that Li measurements are available) and RAVE DR6, to produce a sample of stars with three independent RV determinations. Starting from this sample, we then selected giants following the recipe by Martell et al. (2021) (their Section 2.2), which uses the stellar parameters from GALAH. Then, we took those stars with no issues in their data, reduction, analysis, and iron abundance determination, summarized by the following flags in GALAH DR3: flag_sp == 0, flag_fe_h == 0, extinction parameter E(B-V)¡0.33 (to avoid difficulties with the analysis of spectra of stars with high extinction), and photometric quality flagged as A, B, or C in the WISE $W_{2}$ band (added in the catalog of GALAH) which is later used for classifying the evolutionary phase of the giants (see Buder et al. 2021, for further details of the parameters). Secondly, the sample is restricted to a range of effective temperature $3000$ K $\leq T_{\mathrm{eff}}\leq$ $5730$ K and surface gravity log(g)$\leq 3.2$, which enclose the expected range for red giants in the first-ascendant RGB and the He-core burning phase. Next, we identified if there were any stars belonging to the SMC or LMC by using their spatial and kinematic properties as described by Martell et al. (2021), but no members were found in our sample.
@@
We found a total of $5667$ red giants which have RV measurements with their respective errors reported by the three surveys. We also separately selected those stars that have flag_li_fe == 0 in GALAH DR3, which indicates a line detection good enough to consider a Li abundance measurement without upper limits and flagged problems (Martell et al. 2021; Buder et al. 2021). This subset contains a total of $1418$ red giants. Li abundances were obtained using the logarithmic number density defined as A(X)=$\log(N_{X}/N_{H})+12$ for an element X. From the abundances available in GALAH, i.e., [Li/Fe] and [Fe/H], we computed A(Li)=[[[[Li/Fe]]]]+[[[[Fe/H]]]]+1.05, where A(Li)=1.05 is the abundance value of the Sun from Asplund et al. (2009). The errors for [Li/Fe] and [Fe/H] are given in GALAH DR3, then the errors for the Li abundance are obtained as eA(Li)=$\sqrt{\mathrm{e[Li/Fe]}^{2}+\mathrm{e[Fe/H]}^{2}}$. We found that the mean error is eA(Li)$\sim 0.1$ and the maximum is eA(Li)$\sim 0.25$. Since the A(Li) distributions are later analyzed with a bin size larger than the maximum error (see Section 3.1), we do not apply any additional selection based on eA(Li).
@@ 3.1 RV variability and Li abundance
An important caveat to consider is that some giants can exhibit pulsations, which can produce variations in radial velocities up to $\lesssim 1\ \mathrm{km\ s^{-1}}$ and with typical values ranging from $\sim 10$ to $100\ \mathrm{m\ s^{-1}}$ (e.g. Hekker et al. 2008). We note that $\Delta{RV_{\mathrm{max}}}=359\ \mathrm{m\ s^{-1}}$ is the minimum variation in our giants in classes 5a-5b, with only about $1.1\%$ and $8.4\%$ of them showing variations of $\Delta{RV_{\mathrm{max}}}<500\ \mathrm{m\ s^{-1}}$ and $\Delta{RV_{\mathrm{max}}}<1\ \mathrm{km\ s^{-1}}$, respectively. Thus, while we cannot conclusively distinguish between RV variations caused by pulsations and those induced by companion stars, only a small fraction are likely related to pulsations among our highly variable giants. A combination of both processes can also contribute to RV variations (Beck et al. 2014). Moreover, Cepheids and RRLyrae objects are among the most variable stars due to radial pulsations, with amplitudes $>10\ \mathrm{km\ s^{-1}}$. Given that Gaia DR3 provides a catalog of such objects (Ripepi et al. 2023), we verified that our 1418 giants with measured Li abundances are not part of this catalog.
@@
The right-hand panel of Figure 6 zooms into the high-end tail of the A(Li) distributions from the left-hand panel, starting from A(Li)=1.5. In this panel, we consider joint classes. Thus, $\mathrm{N_{s}/N_{class}}$ is the fraction of stars in a joint class for a given range of A(Li). Here we consider the fractions in the bins normalized with respect to the Li-rich giants in each class. Consequently, given the different normalizations, the shape of the distributions exhibits slight variations compared to what is observed in the left-hand panel of the figure. Giants in class 5a-5b are therefore those Li-rich giants with high certainty of being in a binary system (because they have the largest variation in RV), and they correspond to $37.6\%$ of all the Li-rich stars in our sample. This is only about $5\%$ larger than the fraction of stars with high certainty of variability when considering all giants with A(Li) measurement ($32.9\%$) and all giants with RV in the three surveys ($32.3\%$). Considering error bars, $33\%$ is about one standard deviation from the fraction of Li-rich giants in class 5a-5b (37.6%) which is marginal. Therefore, regardless of their Li abundance, there are about one-third of giants with a high probability of being in a binary system. At the same time, among the Li-rich giants only, there are no significant differences among the A(Li) distributions, they all show that the fraction $\mathrm{N_{s}/N_{class}}$ of giants in each A(Li) bin is almost the same for all the variability classes.